25 bags of wheat flour weighing 20kg costs ₹3250 each. Find the cost if 15 bags of wheat flour, if each bag weighs 8kg

Grade:6

1 Answers

Arun
6 years ago
Dear student
 
25 bags of 20 kg = 500 kg of wheat = 3250 ?
Hence
1 kg of wheat = 3250/500 = 6.5 ?
Hence
15 bags of 8 kg each = 120 kg
Hence 120 kg of wheat = 120 * 6.5 = 780 ?
 
Regards
Arun (askIITians forum expert)